OpenStudy (kiki99):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 123 days and there is an amount equal to 150 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 365 days? Hint: The exponential equation for half-life is A(t) = A0(0.5)t/H, where A(t) is the final amount remaining, A0 is the initial amount, t is time, and H is the half-life.

OpenStudy (kiki99):

A(t) = 150(0.5)t/123, 19.2 grams remaining A(t) = 150(0.5)123t, 1.4 grams remaining A(t) = 123(150)(0.5)t, 0.0 gram remaining A(t) = 150(0.5)123/t, 118.8 grams remaining

OpenStudy (kiki99):

@Kainui

OpenStudy (kiki99):

@RUNNER01

OpenStudy (runner01):

ok lets look

OpenStudy (runner01):

whats the problem

OpenStudy (kiki99):

the statement at the top

OpenStudy (runner01):

i see han on ill give you the answer

OpenStudy (runner01):



OpenStudy (kiki99):

are you sure

OpenStudy (runner01):

i think try and find out if not ill fail this test im on for you

OpenStudy (kiki99):

could you help me with three more questions

OpenStudy (kiki99):

@RUNNER01

OpenStudy (runner01):

pm messages tho ok becuaes i hate when i have people see how i do this

OpenStudy (kiki99):

okay so do i close this question ?

OpenStudy (runner01):

yea